Fermi National Laboratory
 

Operation Summary for November 9, 2011 - November 11, 2011

 

Wednesday November 9

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the MiniBooNE and NuMI experiments taking beam.

During the day shift, a NuMI quadrupole power supply’s (E:QF114) current drifted low and caused the NuMI beam permit to trip off; everything reset. A Booster person completed a study at 9:13 AM. Operators prepared Meson enclosure M03 for access at 9:27 AM for a tour. Operators prepared Meson enclosure M01 for a one-hour access at 9:31 AM to check a sump pump level. Operations disabled the MiniBooNE beam at 9:32 AM in preparation for an access after lunch. MI personnel began an 8 GeV beam study at 9:45 AM. Operations turned the Linac, Booster, MI, and the Transfer Hall at noon in preparation for two hours of accesses and maintenance work. At 12:04 PM, Linac and Booster personnel began RF work. Operators prepared the Muon (formerly Antiproton Source) Prevault and Pretarget areas for access. Operators reported at 3:33 PM that all accesses had ended. Operators had MI ramping by 3:52 PM.

During the evening shift, Operations had established beam to NuMI and MiniBooNE by 4:19 PM. Operations had received permission to established beam to the Meson FTBF experiment T-978 (CALICE) and did so at 6:44 PM. The NuMI beam permit tripped off at 9:01 PM due to bad beam positions; it reset. Operations disabled beam to T-978 at 9:11 PM. Operators reported that a Booster RF station (BRF7) was tripping off so often that the station was bypassed.

Thursday November 10

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE experiments.

During the midnight shift, at 3:49 AM, BRF19 tripped off for the first of three times in about two hours; it reset. At 5:25 AM, the NuMI beam permit tripped off due to bad beam positions; it reset. Operators reported at 5:41 AM that the Tevatron sector A2 sump had been in alarm for some time; operators prepared the area for an access and found the sump pit full of water, the pump on, but not pumping any water out. Operators contacted the Duty Mechanic.

During the day shift, at 8:10 AM, Neutron Therapy Facility (NTF) personnel began preparing to treat one patient later in the morning. Operations established beam to T-978 at 8:12 AM. Safety System personnel reported at 8:26 AM that they were testing the NML ( New Muon Lab) safety system. At 2:47 PM the LRF1 SRM went down, experts found a loose power cable and repaired it. The Muon beam line (formerly PBar) did studies. MI 8.836 Gev studies were also done during the shift.

During the evening shift, Operations did a search and secure for NML. Operations disabled beam to T-978 at 7:02 PM.

Friday November 11

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE experiments.

During the midnight shift, Operations did a search and secure for MT6 Section 2. At 7:14 AM BRF6 tripped off, Operators ran Booster at a lower intensities until experts repaired BRF6. Normal Booster intensities resumed at 7:49 AM.

Operations Coordinator (OpsCo) Plans for Today and the Weekend

Deliver beam to the users.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/11/2011    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory